OpenStudy (anonymous):

PLEASE HELP!!! will give medals!!! 1. which of the following is equal to square root (8x^9y^4 A. 3x^2y^2 B. 3x^4y^2 square root (2x C. 9x^3y^2 D. 3x^3y^2 square root (2 2. what is the simplest form of the expression square root (2)- square root (10)/ square root (2)+square root (10) A. -3/2+1/4 square root (10) B. -3/2+1/2 square root (5) C. 1-1/3 square root (5) D. 1+1/2 square root (5) 3. what is the simplest form of the number 27^2/3 A. 3 B. 6 C. 9 D. 18

OpenStudy (anonymous):

sqrt(8x^9y^4) = sqrt(8)*sqrt(x^9)*sqrt(y^4) = 2sqrt(2)*x^4*sqrt(x)*y^2 = 2x^4y^2 * sqrt(2x)

OpenStudy (anonymous):

okay so wait whats the answer cause i got 3x^y^2 is that right ? for #1

OpenStudy (anonymous):

are you sure that the question is 8x^9y^4 ?

OpenStudy (anonymous):

yes i am it says it on my work

OpenStudy (anonymous):

because the answer is 2x^4y^2 * sqrt(2x) which is none of the options

OpenStudy (anonymous):

it must be 3x^4y^2

OpenStudy (anonymous):

but look at the 3 if you square it you get 9 it is not compatible with the 8 that is given in the question.

OpenStudy (anonymous):

well im just going to put 3x^4y^2 square root (2) do yu know how to do 2 or 3 ? cause im really confused

OpenStudy (anonymous):

this is the question ? (sqrt(3) - sqrt(10)) /(sqrt(2) + sqrt(10))

OpenStudy (charlotte123):

I came here for my shiny medal

OpenStudy (anonymous):

@Coolsector yes it is

OpenStudy (anonymous):

first : 27^2/3 so it is like (27^(1/3) )^2 = 3 ^ 2 = 9

OpenStudy (anonymous):

where did you get 27 ?

OpenStudy (anonymous):

(the third question)

OpenStudy (anonymous):

oh okay i get it

OpenStudy (anonymous):

now for the second: (sqrt(2) - sqrt(10)) /(sqrt(2) + sqrt(10)) multiply and divide by sqrt(2) - sqrt(10) you get: (2 - 2sqrt(2)sqrt(10) + 10) / -8 = -3/2 + (1/4)sqrt(2)sqrt(10) = -3/2 + (1/4) * 2 sqrt(5) = -3/2 + (1/2)sqrt(5)

OpenStudy (anonymous):

oh okay now i get it thank you :)

OpenStudy (anonymous):

yw :)

OpenStudy (anonymous):

still problem with the first though.